Using the table below, calculate the difference in energy between

a) the lowest energy conformations of cis- and trans-1,2-dimethylcyclohexane.

b) the highest and lowest energy conformations of methyl cyclohexane, trans-1,2-dimethylcyclohexane and cis-1,3-dimethylcyclohexane

Type of Strain Interaction Associated Energy
Eclipsing of a pair of H's 4kj/mol
Eclipsing of a H and a CH3 group 6kj/mol
Eclipsing of a pair of CH3 groups 11kj/mol
Gauche (60) Interactin between CH3 Groups 3.8 kj/mol
1,3 CH3 to H Interaction on Cyclohexane Chair 3.8kj/mol
1,3 CH3 to CH3 Interaction on Cyclohexane Chair 15.5kj/mol
Answer #1

1) Difference in lowest energy conformations = 7.6-3.8 = 3.8

2) difference in highest energy conformations =15.2 - 7.6 = 7.6

The lowest energy conformation of methycyclohexane = 0

The highest energy conformation of methylcyclohexane = 7.6 KJ / mole

difference =7.6

b) the energy difference between trans-1,2-dimethylcyclohexane = 15.2-3.8 KJ /mole = 11.4 KJ / mole

c) The energy of lowest conformation of 1,3-dimethylcyclohexane = 0

The energy of highest conformation of 1.3-dimethylcyclohexane = 15.5 KJ / mole